OpenStudy (anonymous):

Answer this for me? <3 A 48 gram sample of a substance that's used for drug research hak-value of 0.1325. N=n0e^-kt N0= initial mass N= mass at time t K= a positice constant that depends on the substance itself and on the units used to measure time t= time, in days Find the substance's half-life, in days. Round answer to the nearest tenth

OpenStudy (amistre64):

answer it for you?

OpenStudy (anonymous):

Please?

OpenStudy (amistre64):

what part of it is giving you troubles?

OpenStudy (amistre64):

if i recall correctly, the time taken to scale it has nothing to do any amounts. ln(scalar)/rate = time

OpenStudy (amistre64):

and the scalar here is 1/2, but id have to dbl chk meself

OpenStudy (amistre64):

N/2 = N e^-kt 1/2 = e^-kt ln(1/2) = -kt ln(1/2)/-k = t

OpenStudy (amistre64):

as i has it it would have just popped out a negative time; to which you can ignore the sign

OpenStudy (amistre64):

all we have to do is plug this into a calculator to find the answer: -ln(1/2)/.1325 = t

OpenStudy (amistre64):

google gives us this: http://www.google.com/search?sourceid=chrome&ie=UTF-8&q=-ln(1%2F2)%2F.1325

OpenStudy (anonymous):

:O , your good !